The length of rectangular label is 2 cm more than the width .If the perimeter of the label is 68 cm find the length and width
Answer by jean14(9)

Let x be the width of rec.
And x + 2cm is the length
With 68cm as Perimeter..we write formula.
68cm = 2( x+ x+2cm) from P= 2 (l x w)
68cm = 2 (2x + 2cm)
68cm = 4x + 4cm
68cm - 4cm = 4x divide both sides by 4
64cm/4 = 4x/4
16cm = x
The width is 16cm.
The length is 16cm +2cm =18cm answer.
Check;
68cm = 2(16cm +18cm)
68cm = 2(34cm)
68cm = 68cm True